OpenStudy (rockindud):

difference and reciprocal of 1(3/4)

OpenStudy (anonymous):

If you actually understood your homework, you'd be able to answer this question easily. I'm sure homework these days have some kind of directions of explanations in them to complete them, read those, and understand them. This isn't a difficult question, good luck.

OpenStudy (rockindud):

say steps to follow plz opigeode i dont want advice

OpenStudy (rockindud):

difference of 1(3/16) and its reciprocal

OpenStudy (anonymous):

Okay, I don't know the difference, but I know the reciprocal, it's 3/16.

OpenStudy (anonymous):

Wait, let me try to find the difference.

OpenStudy (rockindud):

difference of 1(3/16) and its reciprocal options are a) 1(1/8) b)1(1/3) c)15/16 d)105/304

OpenStudy (rockindud):

waiting

OpenStudy (compassionate):

Is this \[1 \times (\frac{ 3 }{ 16})\]

OpenStudy (anonymous):

1 3/4 = 7/4 so the reciprocal is 4/7

OpenStudy (anonymous):

first you have to make into an improper fraction. since it is 1 and 3/4 1 =4 which you add to the top of the fraction. now you get 7/4 the reciprocal is the opposite of the fraction. this means the reciprocal is 4/7